Fourteen years ago, Cathy was twice as old as christian and now the sum of their ages is 40. What is the age of each?
A movie house ticket seller discovered that the end of a day she had $ 2,300 in box. There were two kinds of ticket sold:orchestra tickets costing $6.50 each and balcony tickets costing $10.00 each. The number of orchestra tickets sold were twice as much as the number of tickets sold for the balcony. Find how many tickets were sold for the orchestra and balcony?
i need algebraic soution here

Answer by stanbon(75887)

Fourteen years ago, Cathy was twice as old as christian and now the sum of their ages is 40. What is the age of each?
--------
Now Equation: K + C = 40
Then Equation (K-14)=2(C-14)
K = 2C-14
---------------------------------
Substitute to solve for "C"
2C-14 + C = 40
3C = 54
C = 18 (Christian's age)
---
K + 18 = 40
K = 22 (Cathy's age)
=============================
A movie house ticket seller discovered that the end of a day she had $ 2,300 in box. There were two kinds of ticket sold:orchestra tickets costing $6.50 each and balcony tickets costing $10.00 each. The number of orchestra tickets sold were twice as much as the number of tickets sold for the balcony. Find how many tickets were sold for the orchestra and balcony?
Quantity Equation: R = 2B
Value Equation: 650R + 1000B = 230000 cents
----------------------------
Substitute to solve for "B";
650(2B) + 1000B = 230000
130B + 100B = 23000
230B = 23000
B = 100 (# of balconey tickets sold)
---
R = 2B = 200 (# of orchestra tickets sold)
============================================
Cheers,
Stan H.